DETERMINE THE POINT ON THE GRAPH OF THE EQUATION 2X+5Y = 20 WHOSE X CORDINATE IS 5/2 TIMES ITS ORDINATE .
Answer by lwsshak3(11628)

DETERMINE THE POINT ON THE GRAPH OF THE EQUATION 2X+5Y = 20 WHOSE X CORDINATE IS 5/2 TIMES ITS ORDINATE .
**
2x+5y=20
5y=-2x+20
y=-2x/5+4
This is an equation for of straight line with slope=-2/5,
Which means the x-coordinate is 5/2 times the y-coordinate on all points on the graph in its domain (-∞,∞)